# Anamorphic catadioptric telescope and generalized spherical aberration

### v3

@@ Abstract
Compensation of the generalized spherical aberration components is discussed for the plane-symmetric and anamorphic optical systems. An example compensation rules are derived for the double-plane symmetric optical system containing reflective and refractive subsystems, which focus light in the orthogonal directions at the image plane. An economical anamorphic extension of the Cassegrain-type cylindrical mirror system has been simulated with a ray-tracing code. Anamorphic telescopes of the introduced type provide independent focal distances in the two orthogonal planes allowing to set up a favorable field of view.
@@ 1 Introduction
Anamorphic optical systems (OS), which possess a double-plane symmetry, have been investigated for many years [1, 2, 3, 4, 5, 6, 7] mainly due to their rich capabilities for image formation. However, they reveal also copious structure of aberrations [1, 8, 9, 10, 4, 11, 12, 13, 14] that remain a subject of research. In particular, S. Yuan and J. Sasian have argued that, in the paraxial region, an anamorphic system can be treated as two associated rotationally symmetric OS, and derived for pure mirror systems the generalized ray-tracing equations that include the aberration terms up to the third order [11, 12]. Recently, set of the third-order aberration coefficients in the paraxial approximation was obtained for the systems of cylindrical mirrors or/and lenses directly from the calculation of the optical path difference without approximating these OS by the pairs of rotationally symmetric OS, and as a particular example a cylindrical analog of the Cassegrain telescope was considered [14]. The present research is devoted to its possible anamorphic extensions that allow for two large magnifications in the orthogonal directions at the image plane and good control of the field of view.
@@
We start from discussing the monochromatic third-order aberrations for either anamorphic or plane-symmetric OS, concentrating on the generalized spherical one, in Sect. 2. Next, we introduce a concept of anamorphic catadioptric telescope with two independent focal distances in Sect. 3. We impose compensation of the spherical aberration in the XZ and YZ planes in its initial design and check the performance using a ray-tracing software. We discuss the results and conclude in 4.
2 Plane symmetries and generalized spherical aberrations
@@
The transverse aberrations for paraxial rays propagating in the $Z$ direction through a generic system of cylindrical optical surfaces, which is symmetric under reflection in the $YZ$ plane and translationally invariant along the $Y$ axis, were obtained in Ref. [14] by extension of the classical methods based on direct calculation of the optical path difference [15]. We consider generalization of this system to a single-plane or double-plane symmetric OS. In the later case the $Z$ axis coincides with the crossing line of the two symmetry planes, and the OS may contain anamorphic surfaces, e.g., toroidal ones. The wave aberration function $W$ being the optical path difference between real and intended ideal wavefronts can be expressed in terms of the four transverse coordinates {$x_{0}$, $y_{0}$, $x$, $y$} of the points $P_{0}(x_{0},y_{0},0)$ and $P(x,y,z)$, in which arbitrary ray crosses the object plane and the exit pupil plane, respectively [16]. The both planes are orthogonal to the $Z$ axis. Symmetry in either $XZ$ or $YZ$ plane as well as double-plane symmetry in both these planes leaves the following six second order invariants for the considered four variables
@@
This form contains three constant piston terms (involving only $x_{0}$ and $y_{0}$) and two pairs of identical terms with either $x_{0}^{2}x^{2}$ or $y_{0}^{2}y^{2}$. This leaves in the fourth-order wave aberration function
@@
sixteen different terms that generate third-order aberrations of spherical, comatic, astigmatic (plus field curvature) and distortion types, corresponding to the lines in Eq. (2[17, 18, 19]. For paraxial rays the components of the third-order transverse aberrations can be calculated as
@@
where $\tilde{a}_{1,2}=4a_{1,2}D$, $\tilde{a}_{3}=2a_{3}D$. Clearly, in the case of a single-plane or double-plane symmetry GS aberrations depend on three parameters, which is similar to the case of a single-plane symmetry combined with translational invariance [14]. The GS aberrations degenerate to the ordinary spherical aberration for $\tilde{a}_{1}=\tilde{a}_{2}=\tilde{a}_{3}$. Hence, for the rays in the $XZ$ or $YZ$ planes the coefficients $\tilde{a}_{1}$ and $\tilde{a}_{2}$, respectively, play the role of a usual spherical aberration coefficient.
@@
3 GS aberrations for anamorphic catadioptric telescope
In this section we present an anamorphic telescope [21] with a separate focusing along the orthogonal directions $X$ and $Y$ in case of the optical axis coinciding with the $Z$ axis. To be more specific, we discuss catadioptric schemes consisting of the two sets of optical elements: reflective set (I) and refractive set (II). The set (I) contains cylindrical mirror(s) and is responsible for focusing of the collimated light parallel to the Z axis to a focal segment (in case of absence of the second set) at the focal distance $F_{1}$. The set (II) with smaller focal distance $F_{2}$ contains anamorphic lenses and is responsible for the light pencil compressing in the direction of the focal segment and its focusing to a focal point. The set (II), being an anamorphic extension for the set (I), breaks the translational invariance of the OS to the second plane symmetry.
For definiteness we take the translationally invariant set (I) of Cassegrain type [14], which consists of the two cylindrical mirrors: concave primary M1 and convex secondary M2, as shown in Fig. 2. The focal segment $AB$ of the primary M1 is parallel to the focal segment $A_{2}B_{2}$ of the secondary M2. For the set (II) we concentrate on two economical examples, containing either two or three optical surfaces, which are introduced in the following.
@@ 3.1 Single lens set (II)
We add a toroidal lens L drawn in Fig. 3 behind the mirror set (I) in such a way that its symmetry axis coincides with the focal line $A^{\prime}B^{\prime}$ of the set (I). This lens is a body of rotation of its YZ profile with respect to the axis $A^{\prime}B^{\prime}$. We choose L with a biconvex YZ profile. The OS focal point $C$ is in the center of the segment $A^{\prime}B^{\prime}$. The entire anamorphic OS contains only four optical surfaces, indexed by 1 and 2 for M1 and M2, respectively, and 3 and 4 for the surfaces of L. We describe the cylindrical surfaces by the equation
@@
Figure 3: A schematic isometric view of the optical surfaces 3 (upper) and 4 (lower) of the toroidal lens L.
In the following we discuss aberrations for the considered anamorphic OS, concentrating on GS aberration for brevity. We introduce the conditions for its compensation in the $XZ$ and $YZ$ planes, which correspond to the dotted and dashed lines in Fig. 1, respectively, where the dot-dashed line corresponds to the fulfillment of the both conditions at once. We remark that the introduced method can be applied also to reduction of comatic and other aberration types.
@@ 3.1.1 $XZ$ plane
For paraxial rays in the $XZ$ plane and the marginal ray height at the primary the contribution to GS aberration for the system is determined by the function [14]
@@
where $x_{i}$ is the marginal ray height at the $i$-th surface. The aberration coefficients for the mirrors are given by
@@
To derive the aberration coefficients $A_{313}$ and $A_{314}$ we replace the refractive surfaces 3 and 4 by the cylindrical ones, which are tangential to them in the $XZ$ plane and are described by Eq. (10) with the overall plus sign. Then we use the following generic expression [14]
$A_{31i}=-\,\frac{1}{8}\!\left[\frac{\kappa}{R^{3}}(n_{1}-n_{0})+\frac{n_{0}}{s_{0}}\left(1+\frac{n_{0}}{n_{1}}\frac{s_{0}}{s_{1}}\right)\!\left(\frac{1}{s_{0}}+\frac{1}{R}\right)^{2}\right]_{i}.$ (13)
The $XZ$ projections for the 3rd and 4th surfaces are the circle arcs with the radii $R_{3}$ and $R_{4}$, respectively, which are centered in the OS focal point $C$, see Fig. 2 (left). We take the refractive index to be $n$ inside the lens and unity outside. For $\kappa_{3x}=\kappa_{4x}=0$, where the index $x$ refers to the section by the $XZ$ plane, we obtain
@@
$F_{XZ}\equiv\kappa_{1}+1-\frac{k_{x}^{4}}{\rho^{3}}\left[\kappa_{2}+\left(\frac{\mu+1}{\mu-1}\right)^{2}\right]=0,$ (17)
where $\rho=R_{2}/R_{1}$ and $k_{x}=x_{2}/x_{1}$. In particular, for a telescope of classical Cassegrain type: $\kappa_{1}=-1$ (paraboloidal M1) and $\kappa_{2}=-(\mu+1)^{2}(\mu-1)^{-2}<-1$ (hyperboloidal M2).
@@ 3.1.2 $YZ$ plane
where $y_{i}$ is the marginal ray distance from the $Z$ axis at the $i$-th surface. The aberration coefficients $\tilde{B}_{3i}=A_{32i}$ vanish. To obtain the coefficients $B_{33}^{*}$ and $B_{34}^{*}$ we use the cylindrical surfaces, which are tangential to the surfaces 3 and 4 in the $YZ$ plane and translationally invariant along the $X$ axis, in contrast to the translation along the $Y$ axis in Ref. [14]. To indicate this difference we replace squiggles with stars. This follows in the interchanged roles of $A\leftrightarrow B$ and $x_{i}\leftrightarrow y_{i}$, which results in $B_{3i}^{*}=A_{31i}$ [14]. By labeling curvature radii in the YZ plane with small letters, according to the proper overall sign in Eq. (10), we obtain
@@
where we substitute $s_{03y}=\infty$ and $s_{04y}=s_{13y}-d_{L}$ ($\kappa_{i}=0$ in the paraxial approximation). The thickness $d_{L}$ of of the lens L along the optical axis is given by
$d_{L}=R_{3}-R_{4}.$ (22)
It can not be smaller than the sum of sags $d_{3}$ and $d_{4}$, which can be expressed as
@@
Taking into account that
$\frac{y_{4}}{y_{3}}=1-\frac{n+1}{n}\,\frac{d_{L}}{r_{3}}\approx 1,$ (27)
for $r_{3}=r_{4}\equiv r$ Eq. (18) results in the following condition for compensation of GS aberrations in the $YZ$ plane
$\kappa_{3}+\kappa_{4}=-\frac{1}{n^{2}}-\frac{n}{\nu^{3}}\left(1+n^{2}+n\nu\right)\left(1-n+\nu\right)^{2},$ (28)
where $\nu=n-(n-1)d_{L}r^{-1}$. In Fig. 4 we show this sum of the conic constants as a function of the ratio $d_{L}/r$ for several choices of $n$. Clearly, both $\kappa_{3}$ and $\kappa_{4}$ cannot be zero simultaneously for any reasonable set of parameters, i.e. for not impractically thick lens or high refractive index. Hence, at least one of the lens surfaces must be ‘aspheric’ in its YZ profile to satisfy Eq. (28).
Figure 4: The sum of the conic constants for the surfaces 3 and 4 in the YZ plane versus the ratio of the lens thickness $d_{L}$ to the radius of curvature $r\equiv r_{3}=r_{4}$ for the chosen values of the refractive index $n$: 1.5, 1.7 and 1.9.

### v4

# Two-step focusing anamorphic catadioptric telescope and componentwise aberration compensations

@@ Abstract
Aberration compensation with emphasis on the generalized spherical aberration components is discussed for plane-symmetric and anamorphic optical systems. A narrow field-of-view double-plane symmetric telescope objective containing reflective and refractive subsystems, which focus light separately in the two dimensions of the image plane, is introduced. Analytical derivation of the compensation rules is discussed for the telescope being composed of a Cassegrain-type cylindrical reflective subsystem and an economical refractive one that is determined by up to six essential parameters. Performance of a more advanced version of this telescope, which contains a less aberrative nine-parameter refractive subsystem, is studied using the ray-tracing simulator CODE V. The telescopes of the introduced type provide two independent focal distances for the discussed focusings in the orthogonal dimensions allowing to set up a favorable field of view, and potentially can be used to achieve larger effective aperture area and resolution on a small spacecraft.
@@ 1 Introduction
Anamorphic optical systems (OS), which possess a double-plane symmetry, have been investigated for many years [1, 2, 3, 4, 5, 6, 7] mainly due to their rich capabilities for image formation. However, they reveal also copious structure of aberrations [1, 8, 9, 10, 4, 11, 12, 13, 14] that remain a subject of research. In particular, S. Yuan and J. Sasian have argued that, in the paraxial region, an anamorphic system can be treated as two associated rotationally symmetric OS, and derived the generalized ray-tracing equations that include the aberration terms up to the third order [11, 12]. Recently, set of the third-order aberration coefficients in the paraxial approximation was obtained for the systems of cylindrical mirrors or/and lenses directly from the calculation of the optical path difference without approximating these OS by the pairs of rotationally symmetric OS, and as a particular example a cylindrical analog of the Cassegrain telescope was considered [14]. The present research is devoted to its possible anamorphic extensions. Historically, anamorphic systems have been used in wide-screen imaging since 1950s [4], in laser beam profile controlling since 1970s [2], in semiconductor chip inspection since 1990s [3], etc. However, in this article we are interested in yet another possible type of anamorphic solutions, which allow for two large magnifications in the orthogonal directions at the image plane and good control of the field of view. These solutions can help in fitting a telescope, which collecting area is larger in one dimension, into a space restricted payload or fairing. Rotating such an anamorphic telescope, which provides high spatial resolution in 1D, one can obtain high resolution 2D images after post processing [15]. This gives effectively a larger collection area and resolution as compared to what is achievable with a rotationally symmetric design.
@@
We start from a brief discussion of third-order aberrations for either anamorphic or plane-symmetric OS, concentrating on the generalized spherical aberration, in Sect. 2. Next, we introduce the concept of a catadioptric telescope with two independent focal distances in Sect. 3. We impose spherical aberration compensation in the XZ and YZ planes in its initial design and study performance using a ray-tracing software. We discuss our findings and conclude in Sect. 4.
2 Plane symmetries and aberrations
@@
The transverse primary aberrations for paraxial rays passing in the Z direction through a generic system of cylindrical optical surfaces, which is symmetric under reflection in the YZ plane and translationally invariant along the Y axis, were obtained in Ref. [14] by applying a classical method based on direct calculation of the optical path difference [16]. We consider generalization of this OS to a single-plane or double-plane symmetric one. In the later case the Z axis coincides with the crossing line of the two symmetry planes, and the OS may contain anamorphic surfaces, e.g., toroidal ones. The wave aberration function $W$ being the optical path difference between real and intended ideal wavefronts can be expressed in terms of the four transverse coordinates {$x_{0}$, $y_{0}$, X, Y} of the points $P_{0}(x_{0},y_{0},0)$ and $P(x,y,z)$, in which arbitrary ray crosses the object plane and the exit pupil plane, respectively [17]. The both planes are orthogonal to the Z axis. Symmetry in either XZ or YZ plane as well as double-plane symmetry in both these planes leaves the following six second order invariants for the considered four variables
@@
This form contains three constant piston terms (involving only $x_{0}$ and $y_{0}$) and two pairs of identical terms with either $x_{0}^{2}x^{2}$ or $y_{0}^{2}y^{2}$. This leaves in the fourth-order wave aberration function [18, 19, 20]
@@
sixteen different terms that generate third-order aberrations of spherical, comatic, astigmatic (plus field curvature) and distortion types, corresponding to the four lines in (2). For paraxial rays the components of the third-order transverse aberrations can be calculated as
@@
where $\tilde{a}_{1,2}=4a_{1,2}D$, $\tilde{a}_{3}=2a_{3}D$. Clearly, in the case of a single-plane or double-plane symmetry GS aberrations depend on three parameters, which is similar to the case of a single-plane symmetry combined with translational invariance [14]. The GS aberration degenerates into ordinary spherical aberration for $\tilde{a}_{1}=\tilde{a}_{2}=\tilde{a}_{3}$. Hence, for the rays in the XZ or YZ planes the coefficients $\tilde{a}_{1}$ and $\tilde{a}_{2}$, respectively, play the role of a usual spherical aberration coefficient.
@@
3 Generalized spherical aberration for anamorphic catadioptric telescope
In this section we present an anamorphic telescope [15] with a separate focusing along the orthogonal directions X and Y, and the optical axis coinciding with the Z axis. To be more specific, we discuss catadioptric schemes consisting of the two sets of optical elements: reflective set (I) and refractive set (II). The set (I) contains cylindrical mirror(s) and is responsible for focusing collimated light, which is parallel to the Z axis, to a focal segment (in case of absence of the second set) at the focal distance $F_{1}$. The set (II) with smaller focal distance $F_{2}$ contains anamorphic lenses and is responsible for compressing the light pencil in the direction of the focal segment and its focusing to a focal point. The set (II), being an anamorphic extension for the set (I), breaks the translational invariance of the OS to the second plane symmetry.
For definiteness we take the translationally invariant set (I) of Cassegrain type [14], which consists of the two cylindrical mirrors: the concave primary M1 and the convex secondary M2, as shown in Fig. 2. The focal segment $AB$ of the primary M1 is parallel to the focal segment $A_{2}B_{2}$ of the secondary M2. In our analytical consideration in the next two subsections 3.1 and 3.2 we concentrate on the economical examples of the set (II) containing either two or three optical surfaces, respectively. Then we perform ray-tracing for a more advanced (and less aberrative) version of the discussed telescope with a four-surface set (II) in the subsection 3.3, and discuss its field of view in the subsection 3.4.
@@ 3.1 Single lens set (II)
We add a generalized toroidal shape lens L behind the mirror set (I) in such a way that the symmetry axis of the respective torus coincides with the focal line $A^{\prime}B^{\prime}$ of the set (I). An example lens L is shown in Fig. 3 for the case of minimal thickness (sharp edges) that is not necessary. This lens is a body of rotation of its YZ profile with respect to the axis $A^{\prime}B^{\prime}$. We choose L with a biconvex YZ profile. The OS focal point $C$ is in the center of the segment $A^{\prime}B^{\prime}$. The entire OS contains only four optical surfaces, indexed by 1 and 2 for M1 and M2, respectively, and 3 and 4 for the surfaces of L. We describe the cylindrical surfaces by the equation
@@
Figure 3: A schematic isometric view of the optical surfaces 3 (upper) and 4 (lower) of the toroidal lens L. The lens thickness is discussed in section 3.1.2, its edges are not sharp in practice.
In the following we discuss aberrations for the considered anamorphic OS, concentrating on GS aberration for brevity. We introduce the conditions for compensation of its components in the XZ and YZ planes, which correspond to the dotted and dashed lines in Fig. 1, respectively, where the dot-dashed line corresponds to the fulfillment of the both conditions at once. We remark that the introduced method can be applied also to reduction of comatic and other aberration types.
@@ 3.1.1 XZ plane
Following the method of Ref. [14], we determine contribution to GS aberration for the considered system for paraxial rays in the XZ plane and the marginal ray height at the primary (indicated by the superior ‘(1)’) by means of the following function
@@
where $x_{i}$ is the marginal ray height at the $i$-th surface, the first subindex of the ‘$A$’ quantities refers to spherical aberration ($j=3$ in terms of Refs. [14, 16]), the subindex $s$ goes for (whole optical) system, the second subindex of the aberration coefficients $A_{ijk}$ indicates the aberration component related to the XZ plane, and the third subindex of $A_{ijk}$ numbers the optical surface. The aberration coefficients for the mirrors are given by
@@
To derive the aberration coefficients $A_{313}$ and $A_{314}$ we replace the refractive surfaces 3 and 4 by the cylindrical ones, which are tangential to them in the XZ plane and are described by (10) with the overall plus sign. Then we use the following generic expression [14]
$A_{31i}=-\,\frac{1}{8}\!\left[\frac{\kappa}{R^{3}}(n_{1}-n_{0})+\frac{n_{0}}{s_{0}}\left(1+\frac{n_{0}}{n_{1}}\frac{s_{0}}{s_{1}}\right)\!\left(\frac{1}{s_{0}}+\frac{1}{R}\right)^{2}\right]_{i},$ (13)
where the $i$ index outside the quadratic parenthesis means that all the quantities inside that parenthesis refer to the $i$-th surface. The XZ projections for the 3rd and 4th surfaces are the circle arcs with the radii $R_{3}$ and $R_{4}$, respectively, which are centered in the OS focal point $C$, see Fig. 2 (left). We take the refractive index to be $n$ inside the lens and unity outside. For $\kappa_{3x}=\kappa_{4x}=0$, where the index X refers to the section by the XZ plane, we obtain
@@
$F_{XZ}\equiv\kappa_{1}+1-\frac{k_{x}^{4}}{\rho_{21}^{3}}\left[\kappa_{2}+\left(\frac{\mu+1}{\mu-1}\right)^{2}\right]=0,$ (17)
where $\rho_{21}=R_{2}/R_{1}$ and $k_{x}=x_{2}/x_{1}$. In particular, for a telescope of classical Cassegrain type: $\kappa_{1}=-1$ (parabolic M1 profile) and $\kappa_{2}=-(\mu+1)^{2}(\mu-1)^{-2}<-1$ (hyperbolic M2 profile).
@@ 3.1.2 YZ plane
where $y_{i}$ is the marginal ray distance from the Z axis at the $i$-th surface, the first subindex of all $B$s goes for spherical aberration, while their second subscript either refers to the system ‘$s$’ or numbers the optical surface. The aberration coefficients $\tilde{B}_{3i}\equiv A_{32i}$ vanish. (The coefficients referred to cylindrical surfaces, which YZ profiles are linear, we mark with tildes as in Ref.  [14], and the coefficients referred to generalized toroidal surfaces we mark with stars.) To obtain the coefficients $B_{33}^{*}$ and $B_{34}^{*}$ we use the cylindrical surfaces, which are tangential to the surfaces 3 and 4 in the YZ plane and translationally invariant along the X axis, in contrast to the translation along the Y axis in Ref. [14]. To indicate this difference we replace squiggles with stars. This follows in the interchanged roles of $A\leftrightarrow B$ and $x_{i}\leftrightarrow y_{i}$, which results in $B_{3i}^{*}=A_{31i}$ [14]. By labeling curvature radii in the YZ plane with lowercase letters, according to the proper overall sign in (10), we obtain
@@
where we substitute $s_{03y}=\infty$ and $s_{04y}=s_{13y}-d_{L}$ ($\kappa_{i}=0$ in the paraxial approximation). The central thickness $d_{L}$ of the lens L is given by
$d_{L}=R_{3}-R_{4},$ (22)
where the capital ‘$R$’ denote the radii of curvature in the XZ plane, according to our notation. It can not be smaller than the sum of sags $d_{3}$ and $d_{4}$, which can be expressed as
@@
According to Eqs. (18), (24) and (25), this is convenient to describe the discussed aberration cancellation in terms of the function
$F_{\kappa_{3}\kappa_{4}}=\kappa_{3}+\kappa_{4}\left(\frac{r_{3}}{r_{4}}\right)^{3}\left(\frac{y_{4}}{y_{3}}\right)^{4},$ (27)
which is linear in terms of the conic constants and vanishes along with them, i.e. for spherical lens. Denoting $\rho_{34}\equiv r_{3}/r_{4}$ and taking into account that
$\frac{y_{4}}{y_{3}}=1-\frac{n+1}{n}\,\frac{d_{L}}{r_{3}},$ (28)
we rewrite the condition for GS aberration compensation in the YZ plane as
$F_{\kappa_{3}\kappa_{4}}$ $=-\frac{1}{n^{2}}-\frac{n\rho_{34}^{3}}{\nu^{3}}\left(1+n^{2}+n\nu\right)\left(1-n+\nu\right)^{2}$
×$\times\left(1-\frac{n+1}{n\rho_{34}}\frac{d_{L}}{r_{4}}\right)^{4}.$ (29)
In Fig. 4 we show the dependence of this function on the ratio $d_{L}/r_{4}$ for the case of equal radii $r_{3}=r_{4}$ and several choices of $n$. Then in Figs. 5 and 6 we plot this dependence for the chosen values of the ratio $\rho_{34}$ and the refractive index of 1.5 and 1.9, respectively.
Clearly, the function $F_{\kappa_{3}\kappa_{4}}$ does not vanish, and therefore both $\kappa_{3}$ and $\kappa_{4}$ cannot be zero simultaneously. Hence, at least one of the lens surfaces must be ‘aspheric’ in its YZ profile in order to make it free of the GS aberration component. This result was expected since the mirrors 1 and 2 do not aberrate the rays in the YZ plane, while the considered toroidal lens effects these rays in the same way as a spherically symmetric one does.
Figure 4: The function of conic constants $F_{\kappa_{3}\kappa_{4}}$ versus the ratio of the lens thickness $d_{L}$ to the radius of curvature $r_{4}$ for $r_{3}=r_{4}$ and the chosen values of the refractive index $n$: 1.5, 1.7 and 1.9.
Figure 5: The function of conic constants $F_{\kappa_{3}\kappa_{4}}$ versus the ratio of the lens thickness $d_{L}$ to the radius of curvature $r_{4}$ for the refractive index $n=1.5$ and the chosen values of the ratio of radii $\rho_{34}$: 1, 0.75 and 0.5.
Figure 6: The function of conic constants $F_{\kappa_{3}\kappa_{4}}$ versus the ratio of the lens thickness $d_{L}$ to the radius of curvature $r_{4}$ for the refractive index $n=1.9$ and the chosen values of the ratio of radii $\rho_{34}$: 1, 0.75 and 0.5.